OpenStudy (anonymous):

Find the slope of the line that contains the following pair of points. (4, 3), (2, 8) A. -5/2 B. 2/5 C. 6/11 D. 11/6

OpenStudy (anonymous):

it would be good to have graph paper handy. you might want to graph those points. slope is rise over run. how many spaces to you go up or down? then run is how many spaces you run left or right. hope this helped

OpenStudy (anonymous):

 then plug the points in

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ 3}})\quad &({\color{red}{ 2}}\quad ,&{\color{blue}{ 8}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (anonymous):

slope = (y2 - y1)/(x2-x1) = (8-3)/(2-4) = 5/(-2)

OpenStudy (anonymous):